40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 144 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  9831   Fri Apr 18 19:05:17 2014 jamieUpdateCDSmx_stream not starting on c1ioo

Quote:

To fix open-mx connection to c1ioo, had to restart the mx mapper on fb machine. Command is /opt/mx/sbin/mx_start_mapper, to be run as root. Once this was done, omx_info on c1ioo computer showed fb:0 in the table and mx_stream started back up on its own. 

Thanks so much Rolf (and Keith)!

  9830   Fri Apr 18 14:00:48 2014 rolfUpdateCDSmx_stream not starting on c1ioo

 

 To fix open-mx connection to c1ioo, had to restart the mx mapper on fb machine. Command is /opt/mx/sbin/mx_start_mapper, to be run as root. Once this was done, omx_info on c1ioo computer showed fb:0 in the table and mx_stream started back up on its own. 

  9829   Fri Apr 18 12:53:54 2014 JenneUpdateLSCAttempt at handing CARM to REFL11: Time series

Some time series data from Wednesday night. 

Here is the first time we've gotten the arm transmissions to about 1 count, while holding CARM and DARM on IR signals, so the transmission, as well as POPDC, were relatively quiet.

CARMin1_POPDC_TRX_1000sec_TRXwas1ct.png

As we were attempting to transition CARM to REFL11I, at least 2 of the times, we were hitting a CARM oscillation:

lockloss_CARMoscillating.png

lockloss_CARMoscillating_2.png

  9828   Fri Apr 18 09:40:28 2014 SteveUpdateElectronicsit is a kick

10 days of on-off glitching?   It is a kick. The LSC is off, so it must be the ALS

Attachment 1: theLifeofGlithes.png
theLifeofGlithes.png
Attachment 2: glitchingEndsMC2.png
glitchingEndsMC2.png
Attachment 3: itisaKick.png
itisaKick.png
  9827   Thu Apr 17 17:27:32 2014 ericqUpdateLSCSome reference Plots

Jenne made some suggestions for some plots that would be useful on our CARM offset reduction adventures, so I made some with my MIST model. 

First, here's a plot showing the transfer function of CARM to TRX, with logarithmically spaced offsets out to 3nm. While not a control signal, it shows us where the optical plant resonance stuff is happening. The peaks in this TF correspond to peaks in REFL11, REFL55, AS11, etc., as in the close-to-resonance TFs in ELOG 9785

carm2TRX.pdf

[more to come, had a MATLAB issue]

 

Attachment 2: carmSignalLevels.pdf
carmSignalLevels.pdf
  9826   Thu Apr 17 17:22:32 2014 JenneUpdateCDSmx_stream not starting on c1ioo, locking okay

Jamie tells me that the 2 big consequences of this are (a) we are not archiving any data that is collected on the ioo machine, and (b) that we will not have access to test points on the IOO or ALS models.

To make sure that this is not a show-stopper for locking, I have locked the arms using ALS.  The signals seem to still be getting from the ALS model to the LSC model, and I'm able to acquire ALS lock, so we should be able to work tonight.  All of the data that I have been looking at lately has been coming off of the LSC machine, so we should even be okay in terms of look-back for lockloss studies, etc.

 

  9825   Thu Apr 17 17:15:54 2014 jamieUpdateCDSmx_stream not starting on c1ioo

While trying to get dolphin working on c1ioo, the c1ioo mx_stream processes mysteriously stopped working.  The mx_stream process itself just won't start now.  I have no idea why, or what could have happened to cause this change.  I was working on PCIe dolphin stuff, but have since backed out everything that I had done, and still the c1ioo mx_stream process will not start.

mx_stream relies on the open-mx kernel module, but that appears to be fine:

controls@c1ioo ~ 0$ /opt/open-mx/bin/omx_info  
Open-MX version 1.3.901
 build: root@fb:/root/open-mx-1.3.901 Wed Feb 23 11:13:17 PST 2011

Found 1 boards (32 max) supporting 32 endpoints each:
 c1ioo:0 (board #0 name eth1 addr 00:14:4f:40:64:25)
   managed by driver 'e1000'
   attached to numa node 0

Peer table is ready, mapper is 00:30:48:d6:11:17
================================================
  0) 00:14:4f:40:64:25 c1ioo:0
  1) 00:30:48:d6:11:17 c1iscey:0
  2) 00:25:90:0d:75:bb c1sus:0
  3) 00:30:48:be:11:5d c1iscex:0
  4) 00:30:48:bf:69:4f c1lsc:0
controls@c1ioo ~ 0$ 

However, if trying to start mx_stream now fails:

controls@c1ioo ~ 0$ /opt/rtcds/caltech/c1/target/fb/mx_stream -s c1x03 c1ioo c1als -d fb:0
c1x03
mmapped address is 0x7f885f576000
mapped at 0x7f885f576000
send len = 263596
OMX: Failed to find peer index of board 00:00:00:00:00:00 (Peer Not Found in the Table)
mx_connect failed
controls@c1ioo ~ 1$ 

I'm not quite sure how to interpret this error message.  The "00:00:00:00:00:00" has the form of a 48-bit MAC address that would be used for a hardware identifier, ala the second column of the OMC "peer table" above, although of course all zeros is not an actual address.  So there's some disconnect between mx_stream and the actually omx configuration stuff that's running underneath.

Again, I have no idea what happened.  I spoke to Rolf and he's going to try to help sort this out tomorrow.

Attachment 1: c1ioo_no_mx_stream.png
c1ioo_no_mx_stream.png
  9824   Thu Apr 17 16:59:45 2014 jamieUpdateCDSslightly more successful attempt to get Dolphin working on c1ioo

So it turns out that the card that Rolf had given me was not a Dolphin host adapter after all.  He did have an actual host adapter board on hand, though, and kindly let us take it.  And this one works!

I installed the new board in c1ioo, and it recognized it.  Upon boot, the dolphin configuration scripts managed to automatically recognize the card, load the necessary kernel modules, and configure it.  I'll describe below how I got everything working.

However, at some point mx_stream stopped working on c1ioo.  I have no idea why, and it shouldn't be related to any of this dolphin stuff at all.  But given that mx_stream stopped working at the same time the dolphin stuff started working, I didn't take any chances and completely backed out all the dolphin stuff on c1ioo, including removing the dolphin host adapter from the chassis all together.  Unfortunately that didn't fix any of the mx_stream issues, so mx_stream continues to not work on c1ioo.  I'll follow up in a separate post about that.  In the meantime, here's what I did to get dolphin working on c1ioo:

c1ioo Dolphin configuration

To get the new host recognized on the Dolphin network, I had to make a couple of changes to the dolphin manager setup on fb.  I referenced the following page:

https://cdswiki.ligo-la.caltech.edu/foswiki/bin/view/CDS/DolphinHowTo

Below are the two patches I made to the dolphin ("dis") config files on fb:

--- /etc/dis/dishosts.conf.bak    2014-04-17 09:31:08.000000000 -0700
+++ /etc/dis/dishosts.conf    2014-04-17 09:28:27.000000000 -0700
@@ -26,6 +26,8 @@
 ADAPTER:  c1sus_a0 8 0 4
 HOSTNAME: c1lsc
 ADAPTER:  c1lsc_a0 12 0 4
+HOSTNAME: c1ioo
+ADAPTER:  c1ioo_a0 16 0 4
 
 # Here we define a socket adapter in single mode.
 #SOCKETADAPTER: sockad_0 SINGLE 0

--- /etc/dis/networkmanager.conf.bak    2014-04-17 09:30:40.000000000 -0700
+++ /etc/dis/networkmanager.conf    2014-04-17 09:30:48.000000000 -0700
@@ -39,7 +39,7 @@
 # Number of nodes in X Dimension. If you are using a single ring, please
 # specify number of nodes in ring.
 
--dimensionX 2;
+-dimensionX 3;
 
 # Number of nodes in Y Dimension.

I then had to restart the DIS network manager to see these changes take affect:

$ sudo /etc/init.d/dis_networkmgr restart

I then rebooted c1ioo one more time, after which c1ioo showed up in the dxadmin GUI.

At this point I tried adding a dolphin IPC connection between c1als and c1lsc to see if it worked.  Unfortunately everything crashed every time I tried to run the models (including models on other machines!).  The problem was that I had forgotten to tell the c1ioo IOP (c1x03) to use PCIe RFM (i.e. Dolphin).  This is done by adding the following flag to the cdsParamters block in the IOP:

pciRfm=1

Once this was added, and the IOP was rebuilt/installed/restarted and came back up fine.  The c1als model with the dolphin output also came up fine.

However, at this point I ran into the c1ioo mx_stream problem and started backing everything out.

 

  9823   Thu Apr 17 16:04:40 2014 JenneUpdateElectronicsHigh gain Trans PD electronics change

 

 I have made the same modification to the Yarm trans PD whitening board as was done for the xend, to increase our SNR.  I put in a 2.1kOhm thin film resistor in the Rgain place.

When I was pulling the board, the ribbon cable that goes to the ADC had its connector break.  I redid the ribbon connector before putting the board back. 

I see signals coming into the digital system for both the high gain and low gain Y transmission PDs, so I think we're back.  I will re-do the normalization after Jamie is finished working on the computers for the day.

  9822   Thu Apr 17 11:00:54 2014 jamieUpdateCDSfailed attempt to get Dolphin working on c1ioo

I've been trying to get c1ioo on the Dolphin network, but have not yet been successful.

Background: if we can put the c1ioo machine on the fast Dolphin IPC network, we can essentially eliminate latencies between the c1als model and the c1lsc model, which are currently connected via a rube goldberg-esq c1lsc->dolphin->c1sus->rfm->c1ioo configuration.

Rolf gave us a Dolpin host adapter card, and we purchased a Dolphin fiber cable to run from the 1X2 rack to the 1X4 rack where the Dolphin switch is.

Yesterday I installed the dolphin card into c1ioo.  Unfortunately, c1ioo, which is Sun Fire X4600, and therefore different than the rest of the front end machines, doesn't seem to be recognizing the card.  The /etc/dolphin_present.sh script, which is supposed to detect the presence of the card by grep'ing for the string 'Stargen' in the lspci output, returns null.

I've tried moving the card to different PCIe slots, as well as swapping it out with another Dolphin host adapter that we have.  Neither worked.

I looked at the Dolphin host adapter installed in c1lsc and it's quite different, presumably a newer or older model.  Not sure if that has anything to do with anything.

I'm contacting Rolf to see if he has any other ideas.

  9821   Thu Apr 17 01:18:34 2014 JenneUpdateLSCAttempt at handing CARM to REFL11

[Jenne, Koji]

This evening, Koji and I followed the procedure from last night (elog 9817), with the exceptions that as we saw gain peaking in the DARM spectrum, we lowered the DARM servo gain.  We also left the DARM FM4 boost off, and added (TRX+TRY) power normalization to AS55 (although we still had to hand-reduce the gain).    These changes enabled us to reduce the CARM offset much further.  We were able to get the transmitted powers to hold steady at about 1 count while on the IR signals, which is a new record for us.  (We had in the past held the arms with ALS at several counts, but the power fluctuations were huge, and now they are nice and small).

After that, we started looking at whether we could transition CARM over to REFL11I.  We tried a few times, but never made it all the way.


Here are some times for data-foraging tomorrow:

8:27pm, nice transition, CARM offset reduction to 0.6 before lockloss.

9:19pm, turned on power normalization for AS55Q, then reduced CARM offset to 0.5

9:40pm, Lockloss after reducing CARM offset to -0.24, arm transmitted powers around 0.9.

gps 1081748419: First trial trying to transition CARM to REFL11I normalized by (TRX+TRY).

gps 1081749965:  Tried to transition CARM to (REFL11 + REFL33)/(TRX+TRY).  Got about 1/3 of the way through the transition (in terms of matrix element value steps) before lockloss.

11:56pm, Tried to add in REFL11I to CARM error signal (without reducing 1/sqrt(trans) matrix elements).  We increased the REFL11 matrix element until we saw gain peaking, and then tried reducing the 1/sqrt(trans) contribution, and lost lock.  We were only at an offset of 0.3, so we probably weren't close enough to the resonance yet.  We were able to add in REFL11 information, but this was probably not too hard, since there wasn't much actual information in it.


Thoughts:

* It's a little weird that once we are on IR signals, the 0 CARM offset point that we find with ALS is not the true CARM offset point.  Although, this may be because we're just going to an averaged no CARM offset place with ALS, but since ALS is noisy, we won't ever really be holding on the zero offset point.  Anyhow, when we were using the 1/sqrt(trans) signals for CARM, and the CARM digital offset was -0.24, the ALSX and ALSY outputs were both about 0.5 in magnitude.

* We're getting there! 

  9820   Thu Apr 17 01:01:02 2014 JenneUpdateLSCLSC model modifications

Last night, EricQ and I were concerned that we might need some CARM UGF servoing, so I added a UGF servo block, copied from the aLIGO LSC model, to our LSC model.  The block is inline with the CARM servo, after the output triggering, just before the output matrix.  Q put together some screens, which are accessible from the main LSC screen. 

The model is compiled and running.  We didn't get very far in testing it though before Koji pointed out that it is a slow solution, and not a fast one like we were searching for.  We were hoping to deal with the momentary power buildup, and thus optical gain change, as the arms flash close to resonance.  The UGF servo will not work nearly that fast though.  We may want it for slow UGF servo-ing, but it's not the solution to what Q and I were thinking about yesterday.  Regular ol' dynamic normalization is closer to the right answer for this.

In tonight's activities, Koji and I found that we probably want a CESAR block for DARM as well as CARM, so that we can independently normalize AS55Q. 

To solve the DARM oscillation issue from last night (that I discovered this evening when I finally looked at the time series data), we may want to implement a DARM UGF servo.  For tonight, as we reduced the CARM offset and started seeing gain peaking in the DARM spectra, I hand-reduced the DARM gain.

 

  9819   Thu Apr 17 00:49:06 2014 JenneUpdateLSCCARM and DARM on IR signals, boosts engaged

I looked at 2 of the locklosses from last night, (1:19am and 1:27am), and saw that for both, the DARM loop started to oscillate just before we lost lock.  In the trials tonight, we were more watchful of gain peaking.

Here is the 1:19am lockloss

Lockloss_DARMgainTooHigh_119am.png

And here is the 1:27am lockloss

Lockloss_DARMgainTooHigh_127am.png


 So you can see what we were doing, and what the effect was, here is a few minutes of data just before the 1:27am lockloss. The times I note below are rough, but should give you an idea of what happened at which time.

0 sec:  Arms are held on resonance with ALS.

10 sec:  CARM offset of 3nm added.

20 sec:  PRM restored, one flash, then PRMI acquires lock.

30 sec:  CARM offset reduced to 2nm, transmitted powers are about 0.1

50 sec:  Transition CARM to 1/sqrt(trans) signals.  Note that we are using the high gain Thorlabs PD here, so the noise is better than last Thursday.

60-110 sec:  CARM offset reduction to about 1nm.

110 sec:  CARM's LSC low frequency boost engaged.

120 sec:  DARM transitioned to AS55Q.

170 sec:  DARM's LSC low frequency boost engaged.

SmoothCARMandDARMtransitions_LSCboosts.png

  9818   Wed Apr 16 02:29:30 2014 ericqUpdateLSCCARM and DARM on IR signals, boosts engaged

 As Jenne mentioned, we took OLTF transfer functions, and determined that we had more than enough phase margin to switch on the LSC boosts in FM4. This improved the error signal noise spectra quite a lot, and noticeably reduced the TRX/TRY fluctuations, and actuation output. 

Here's the CARM OLTF (FM4 boost on in red, boost off in black)

carmOLTF.pdf

 

Here's what happened to the CARM and DARM spectra when we turned on the boosts. (ALS only in black, initial IR signal transitions in mid-color, boosted IR signals in bright color)

boostPlot.png 

  9817   Wed Apr 16 02:11:40 2014 JenneUpdateLSCCARM and DARM on IR signals, boosts engaged

[Jenne, EricQ]

Tonight, we transitioned CARM and DARM to IR signals, took loop transfer functions, and determined that we could engage the LSC boosts (FM4 in the CARM and DARM servos, which are the same as the XARM and YARM servos). 

Q is preparing spectra to post, and I will dig out time series.  Look for these tomorrow, if they aren't posted tonight.

For the time series data fetching, I have taken notes on what we were doing when, so that I can actually find the data.


11:09pm:  CARM's LSC boost on for the first time

11:14pm:  DARM transferred to AS55Q

11:21pm:  DARM's LSC boost on for the first time

(lockloss)

11:53pm:  CARM transition

12:02am:  DARM transition done, both LSC boosts on

12:04am:  lockloss after reducing CARM digital offset to 0.4

12:45am: PRMI + 2 arms flashing, with no CARM or DARM offsets (arms still on ALS) because we forgot to put in the CARM offset before restoring PRM alignment.  PRMI may have been actually locked, or we may just have been flashing....need to look through the data to see what our recycling looked like.

(lockloss)

1:05am:  pretty smooth transition completed (both CARM and DARM), but we lost lock while reducing the CARM offset.

1:19am: lockloss - why?? We were just sitting at a CARM offset of about 1.3nm (1.3 counts), holding on IR signals.  We were not touching any IFO things while looking at some plots, and just lost lock.  Want to see if we can understand why.

1:27am:  another nice smooth transition for both CARM and DARM to IR signals, but almost immediate lockloss when reducing the CARM offset.


Using the new ALS lock acquisition scripts (elog 9816) and our transition scripts, getting back to PRFPMI lock is pretty smooth and procedural.

* Align arms using ASS (ifo configure screen, restore xarm and yarm, run both arms' ass scripts).

* Align PRMI, no arms (ifo configure screen, restore prmi sideband)

* Find ALS beatnotes, with arm lasers on opposite sides of the PSL.  For both, when increasing the value of the temperature slider, the beatnote should increase in frequency.  (ifo configure screen, restore CARM and DARM als)

* Run ...../scripts/ALS/Lock_ALS_CARM_and_DARM.py

* Run "Find resonance" scripts from ALS screen for each arm.

* Put in a 3 count offset to CARM loop.

* Restore PRM alignment.  (PRMI should acquire lock immediately, although PRM may need some small alignment tweaking).  Enable PRCL and MICH outputs, PRM and BS actuation outputs.

* Reduce CARM offset to 2 counts. 

* Set offsets of 1/sqrt(TRX) and 1/sqrt(TRY) filter banks in the AUXERR section of the LSC screen.  The outputs of both should equal 2 counts (to match the 2 count offset in the CARM loop). 

* Run .../scripts/PRFPMI/Transition_CARM_ALS_to_TransSqrtInv.py , making sure to reduce the CARM digital offset if needed, to keep the arm transmissions at about 0.1 counts.

* Engage FM4 of the CARM filter bank, which is the LSC boost.

* Run .../scripts/PRFPMI/Transition_DARM_ALS_to_AS55.py , making sure to reduce the CARM (or should be DARM?) digital offset if needed, to keep the arm transmissions at about 0.1 counts.

* Engage FM4 of the DARM filter bank, which is the LSC boost.


Notes for going forward:

When we have small-ish digital CARM offsets, such that both of our arm transmitted powers are about 0.1 or higher, we see clear coherence between our CARM_IN1 (the 1/sqrt(trans) signals) and a normalized REFL11_I (again using a spare filter bank like XARM to get REFL11 normalized by (TRX+TRY) ).  We have not yet tried transitioning the CARM digital error signal to this normalized REFL11.

Even though we see that the IFO is much less noisy (as measured by significantly reduced RIN in TRX and TRY as visible by eye on Dataveiwer), we are still losing lock when we reduce the CARM offset.  I have noted above several times, for when we had locklosses, so that I can see if I see anything elucidating in the time series data.

  9816   Wed Apr 16 01:51:16 2014 JenneUpdateLSCScripts written for ALS acquisition, CARM and DARM transitions

[Jenne, EricQ]

This evening, as part of locking activities, we threw together some handy scripts.


The first one, "Lock_ALS_CARM_and_DARM.py" (no judging of my naming style!!), lives in .../scripts/ALS/ . 

It acquires ALS lock in CARM and DARM mode, so we don't have to do it by hand anymore.

The first thing that it does is ask you to acknowledge that your beatnotes are in place, and they follow our new (newer than the last elog about conventions) beatnote convention.  You are reminded in the terminal window what that convention is:  When the temperature sliders for either arm is INCREASED, the beatnote frequency should INCREASE. 

After you acknowledge that the beatnotes are good, it sets the CARM and DARM servo gains to zero, enables the outputs, sets the input matrix elements, clears the phase tracker histories, and starts ramping up the gains (with +1,+1 for DARM, the darm servo gain is +positive.  with -1*ALSX,+1*ALSY for CARM, the carm servo gain is -negative).  At a gain of 3, it engages the integrators and the resonant gains.  At the final gain of 6, it engages the boosts.

We have used this script ~10 times tonight, and it's been great every time.


The next two scripts are for making the transition from ALS to IR signals.  They both live in ..../scripts/PRFPMI/

"Transition_CARM_ALS_to_TransSqrtInv.py" (again - no judging!) slowly blends the input matrix elements to swap CARM control from the ALS signals to the 1/sqrt(trans) signals.  It takes a few steps, and asks for a keyboard input between steps.  This is because if our 1/sqrt(trans) offsets aren't perfect, we can start to lose transmission power.  To mitigate this, we decrease the offset in the CARM servo filter bank to get more power back.  This script requires an input, which is what you want the final sqrtinv matrix elements to be.  It will fail without this.  For a CARM offset, both of the final sqrtinv matrix elements will have the same sign.

"Transition_DARM_ALS_to_AS55.py" (I can telepathically hear you judging me right now.)  does the same blending, except to swap DARM control from ALS signals to AS55Q.  For the same reason of imperfect offset-setting, it takes several steps, to allow you to adjust the CARM offset if needed. Although, after typing this, I realized that perhaps we should have been tweaking the DARM offset.  Either way, this transition required much less tweaking of offsets than the CARM transition did.  Again, the script requires an input, which is your final desired AS55Q->DARM matrix element value.

Both of these scripts should be run at a digital CARM offset of about 2 counts, although with the offset tweaking during the CARM transition, I usually end at about 1.5 counts. 

*  To determine the final gain value for the CARM sqrtinv matrix elements, we have been using a spare filter bank (ex. XARM), and having the input to that be the sum of the sqrtinv channels.  We then put in a CARM line, and look at the transfer function between the temporary filter bank's input, and the CARM_IN1. 

*  To determine the final gain value for the DARM AS55 matrix element, we have been doing a similar thing, looking at the transfer function between DARM_IN1 and AS55Q with a DARM line on.  We have been putting this DC gain into the static PD normalization (4th block from the left on the big LSC screen), although with the new script, it will be easier to just put that value into the matrix element.

  9815   Tue Apr 15 16:21:18 2014 KojiUpdateIOOMC2 LSC offset was set to be -5000

Yesterday, MC2 alignment was slipping all day. Even when the WFS was off (i.e. there wa sno actuation), I had continual misalignment caused by MC2

I was afraid that the MC2 mirror is on a bistable position somehow. So I gave -5000 offset on the MC2 LSC. We'll see how it makes the MC happier.

  9814   Tue Apr 15 13:24:42 2014 SteveUpdatePSLlaser turned on

The 2W Innolight was off for 4 hours.

Attachment 1: laserOff4hrs.png
laserOff4hrs.png
Attachment 2: laserShutsDown.png
laserShutsDown.png
  9813   Tue Apr 15 09:32:19 2014 GabrieleUpdateLSCMICH gets noisy as CARM or DARM offset reduced

I guess this is normal. DARM has (almost) the same effect of MICH on the corner interferometer signals, just increased in amplitude by the arm cavity amplification. When the arm is out of resonance, DARM is almost completely depressed and it is not affecting MICH at all. On the other hand, when the cavities are exactly at resonance, DARM signal is amplified w.r.t. MICH by the cavity gain (2F/pi).

Since DARM is still controlled with ALS, it is probably noisy. The closer to resonance you move the cavities, the more ALS noise in DARM will affect MICH.


Quote:

When looking at the data, I see that the MICH error signal gets fuzzier when the arms get close to resonance. (Note here that because I forgot to zero the carm offset before finding the resonances, -3 is my zero point for this plot and the next.) 

  9812   Tue Apr 15 08:55:57 2014 KojiUpdateLSCAnalog phasing of REFL11 and REFL55

I have never used such a long cable for RF phase adjustment. The speed of the signal is 2e8 m/s and the frequency is ~10e6 Hz.
This means that the wavelength is only about 20m. How could you end up with ~100meters?
The convenient way to remember the cable delay is "1m, 1MHz, 2deg". This gives us ~1.5m for 11MHz and 34deg.

In fact, 1 degree of phase shift is not 1/(2 pi freq) second of delay, but f/360.

For such a precise phase adjustment, it is better to calibrate the delay with the network analyzer.

Quote:

We calculated that about 1 degree of phase shift is about 1/(2 * pi * freq), or about 1.4e-8 seconds of delay for 11MHz.  We took the speed of light in the cables to be about 2/3*c, so 1.4e-8 * 2e8 = 2.9 meters per degree for 11MHz.  Since REFL11 was 34 degrees from 0, we estimate that we need to add about 98 meters of cable to the REFL11 signal path.  The same calculation for 55 MHz, but with a 15 degree shift required, gives 8.8 meters of cable to be added to the REFL55 signal path.   

 

  9811   Tue Apr 15 02:26:45 2014 ericqUpdateLSCAnalog phasing of REFL11 and REFL55

For future reference:

As we were poking around with the common mode servo in an FPMI configuration, we locked CARM/DARM with ALS as in recent ELOGs.

MICH was locked on ASDC: ASDC -> MICH = 10.0 in the DCPD DoF Matrix (I couldn't easily get AS55Q working, ASDC worked quickly and good enough)

MICH gain +25, FM4 FM5 On, FM2 switched on once locked. Offset was manually adjusted to get closer to dark fringe.

Actuated on BS: MICH->BS = 0.5 in Output Matrix.

  9810   Tue Apr 15 02:19:54 2014 JenneUpdateLSCAnalog phasing of REFL11 and REFL55

[Jenne, EricQ]

I told Koji that I wanted to play with the common mode servo this evening, and he pointed out that we only get the signals after the digital demod phase angle in the digital system (obviously).  So, if I want to use either REFL11 or REFL55 for my CARM signal, I want to do something in analog-land so that my digital demod phase is close to 0 or 90. 

While we had the PRFPMI locked (with CARM offset of 2 or 3 nm), we set the demod phases of REFL11 and REFL55 to minimize a CARM line in the Q-phase.  This gave us -34 degrees for REFL11, and -75 degrees for REFL55. 

We calculated that about 1 degree of phase shift is about 1/(2 * pi * freq), or about 1.4e-8 seconds of delay for 11MHz.  We took the speed of light in the cables to be about 2/3*c, so 1.4e-8 * 2e8 = 2.9 meters per degree for 11MHz.  Since REFL11 was 34 degrees from 0, we estimate that we need to add about 98 meters of cable to the REFL11 signal path.  The same calculation for 55 MHz, but with a 15 degree shift required, gives 8.8 meters of cable to be added to the REFL55 signal path. 

I connected up some long BNC cables, and inserted them between the heliax breakout board on the LSC rack, and the respective PD inputs of the REFL11 and REFL55 demod boards.  I used (45 meters + 45 meters + a little bit) for REFL11, and used about 9 meters for REFL55. 

When we relocked the PRFPMI, and redid the phasing, we were very close to zero for both REFL11 and REFL55!  REFL11's digital demod phase is now +1 degree, and REFL55's digital demod phase is -5 degrees.

We changed the input of the CM servo board from POY (which Den and Koji had been using in December - see elog 9500) to REFL11 I MON. 


Q locked the FPMI (separate reply elog), and then we tried engaging the CM analog servo.  We were not successful. 

 

These settings were mostly copied from elog 9500, so they are almost surely not correct. 

CM servo screen:  In1 gain = 31dB, switch on, offset = -2.7V, boost off, super boosts off, option=disable, 79:1.6k switch disabled, polarity minus, option disable, AO gain=8dB, limiter enable.

For the slow path, CM_SLOW -> MC LSC servo had a +1 in the input matrix. 

CM filters in the AUX_ERR screen:  FM1 (unwhite) on, all others off, gain = 2.6. 

MC servo filters:  FM7, FM10 on, all others off (no triggered filter modules).  Gain = 0 initially.

MC servo board AO path disabled initially, G=-32dB initially.

 

Once Q had the FPMI locked, I tried increasing just the CM analog gain (by enabling the AO path on the MC board, and increasing the gain).  Doing this, I lost lock at -3 dB. 

I then tried again, this time alternating increasing the analog gain, and increasing the MC LSC servo gain.  I got up to 3e-3 for the MC digital gain, and -7 dB for the analog gain before we lost lock again.

 

We have determined that we should probably try just locking one of the arms with POX or POY, as Den and Koji did, to get a feel for how the system works.

 

 

  9809   Mon Apr 14 19:02:09 2014 JenneUpdateLSCMICH gets noisy as CARM or DARM offset reduced

This afternoon, I was toying around with reducing either the CARM or DARM offsets (so, put in a CARM offset, leave DARM zero, lock the PRMI, then reduce CARM offset to zero.  Or, put in a DARM offset, leaving CARM offset zero, lock the PRMI, then reduce the DARM offset to zero).

When looking at the data, I see that the MICH error signal gets fuzzier when the arms get close to resonance. (Note here that because I forgot to zero the carm offset before finding the resonances, -3 is my zero point for this plot and the next.)

MICH_fuzzy_when_offsets_small_longerData.png

Here is a zoom of the last piece of this time series, but with both TRX and TRY plotted (along with POPDC, CARM_ERR and DARM_ERR), where you can see that I had a momentary power buildup of > 100 transmission counts, which is about 20% of our final expected power.

TRX_TRY_100cts.png

Here is a different time series, showing a reduction of the DARM offset, and you can see that as the offset approaches zero, the MICH error signal gets noticeably more fuzzy.  Somewhere near the 240 second mark, I lose PRMI lock.

MICH_fuzzy_when_offsets_small.png

  9808   Mon Apr 14 17:59:05 2014 JenneConfigurationPEMNew T-240 cable

As payment for borrowing 2 of our seismometers, Zach has made us a new Trillium cable, to go from the granite station to the readout box, which we can put into 1X7, where the PEM ADC is.  To put the T-240 in side the can, and seal it, we need a little jumper cable from the seismometer to the granite, but for now, we can just pass this cable underneath the can.

  9807   Mon Apr 14 13:20:45 2014 JenneUpdateLSCIFO Configure screen updated, CARM / DARM scripts added

I have compressed the IFO Configure screen.  All PRMI things (sideband lock and carrier lock) are in the PRMI button, all arm things (both RF and ALS) are in the respective arm buttons.

I have also made a new set of scripts for CARM and DARM lock acquisition with ALS. 

I hope that each button's purpose is clear, but take a second to look at them before you next use the IFO Configure screen.

  9806   Mon Apr 14 11:19:55 2014 JenneUpdateLSCMC WFS found off

I'm not sure why, but the WFS were turned off when I came in this morning.  The MC was not staying locked, and even during brief locks, the FSS FAST out was railed at 10. 

Aligning the MC mirrors to maximize the transmission, and then engaging the WFS seems to have made things better.

  9805   Sun Apr 13 13:03:34 2014 ranaUpdateLSC 

That's a very smooth DARM transition - its good news that the dALS signals don't have a huge offset w.r.t the real error signal.

It would be interesting to see if the MICH can be locked and stay locked which CARM is ramped in. We would want to hold it with the Q phase of the CARM PD once its on.

May not be a milestone, but its cool anyway. 

+2 points.

Will also be cool to see how soon the CM servo can be switched on in the acquisition sequence. Maybe ALS_COMM -> CM board, gets mixed with TRXY for low frequencies in the intermediate stage before final RF?

  9804   Fri Apr 11 18:55:28 2014 manasaUpdateLSCarm length measurements

Arm lengths were measured using ALS

X arm length = 37.79 +/- 0.05 m

Y arm length = 37.81 +/- 0.01 m

Whats and whys

We want to measure the arm length with an accuracy of say a mm.

This would mean a measurement precision of 1e-3/40=25ppm. (1mm in 40m)

So the required measurement resolution on the spectrum analyser is 25ppm*4MHz=100Hz (assuming the cavity FSR is roughly 4MHz). 

Although the spectrum analyser does not limit the measurement precision, we are limited by the noise in ALS at 1000Hz rms. So we can use ALS only to measure arm length precise to the order of a few mm.

RXA: Not that we really need to right now, but even with an ALS noise of 1000 Hz, we can can do better just by averaging at each resonance point. And fitting a line as you have already done gets even better:

http://en.wikipedia.org/wiki/Propagation_of_uncertainty

 

The Spectrum analyser was reference locked to the rubidium clock @10MHz for these measurements.

The FSRs of the arms

X arm = 3.9671e+06 +/- 4.8535e+03 Hz

Y arm = 3.9648e+06 +/- 1.1064e+03 Hz

Attachments:

1&2. Plots representing the arm scans showing the beat frequency for which IR resonates in the arm vs the ALS offset (position of the ETM).

3. Data and code (zip file)

P.S. We had trouble scanning the arms using ALS. This was because the slow servo was not enabled. Hence ALS was losing its PDH lock everytime we scanned past a couple of FSRs.

Attachment 1: Xarm.png
Xarm.png
Attachment 2: Yarm.png
Yarm.png
Attachment 3: 40mCavLength.zip
  9803   Fri Apr 11 16:04:31 2014 KojiUpdateLSCcongratulation

It's just one of the stepping stones, but not yet a mile stone.
Keep going forward!

  9802   Fri Apr 11 14:57:53 2014 steveUpdateLSCcongratulation

 

 

Attachment 1: whenSheisHappy3h.png
whenSheisHappy3h.png
Attachment 2: mc4hArms3hrs.png
mc4hArms3hrs.png
  9801   Fri Apr 11 12:32:33 2014 ericqUpdateLSCCARM and DARM both on IR signals!!!!!!!!!

Quote:

How much was the whitening gain for AS55 this time? 

 21 dB. We played with the whitening gain a little bit; at around 30dB with the signal levels at TRX = .1ish, we were consistently saturating the ADC. 

  9800   Fri Apr 11 12:21:27 2014 KojiUpdateLSCCARM and DARM both on IR signals!!!!!!!!!

About the ADC range,

According to the elogs, DARM = AS55Q/400. So in the current level, the error has +/-40cntpp (even if I ignore the whitening).

The arm transmission this time was 0.1-0.3. This will go up to 100~300. So we potentially increase the AS55Q optical gain by factor of 1000.

So we get +/-40000. This is already too much. If we consider the whitening, the situation is more tough.

We need to lower the whitening gain. If it is not enough, we need to lower the power on the PD.

How much was the whitening gain for AS55 this time?

Quote:

 DARM_IN1_getsSmaller.png

 

  9799   Fri Apr 11 11:58:24 2014 JenneUpdateLSCCARM and DARM both on IR signals!!!!!!!!!

 A few time series from last night's data. 

300 seconds, starting from 1081240100, showing that as we move from ALSY-ALSX to AS55Q, the DARM error signal gets smaller.

DARM_IN1_getsSmaller.png

The same 300 seconds, showing that the CARM error signal, and the arm transmissions, are not perturbed during this transition.

DARM_CARM_TRX_TRY.png

DARM in and out, for 300 seconds, showing that the control output also gets smaller.

DARMin_DARMout.png

A slightly longer time series, ending at about the same time, but starting a few minutes earlier, showing us (1) adding a 3 count CARM offset, (2) locking the PRMI (3) transitioning CARM to sqrtinv signals, and then (4) transitioning DARM to AS55Q.

CARMtransition_DARMtransition.png

CARM and DARM in and outs, for the 500 second time chunk showing all the transitions.  Unfortunately, it looks like CARM_OUT is more noisy when it's on the sqrtinv signals, than it was on the ALS signals.  Part of this may be that we have not yet swapped the resistor in the TRY QPD, to improve the SNR in the same way that we have already done for the TRX QPD.  [EDIT, JCD:  Also, we had hard-triggered the Trans switching, so we were only looking at the QPD sum for the TRX and TRY, and the QPDs only have a few ADC counts at low transmissions, so we had poor SNR for that reason too.]

CARM_DARM_in_out.png

  9798   Fri Apr 11 10:30:48 2014 jamieUpdateLSCCARM and DARM both on IR signals!!!!!!!!!

Quote:

[EricQ, Jenne]

We're still working, but I'm really excited, so here's our news:  We are currently holding the IFO on all IR signalsNo green, no ALS is being used at all!!!!  

 Phenomenal!!  Well done, guys!

  9797   Fri Apr 11 02:09:31 2014 JenneUpdateLSCCARM and DARM both on IR signals!!!!!!!!!

[EricQ, Jenne]

A few more details on our work for the evening, of switching the PRFPMI completely to IR signals (although still with a pretty big CARM offset).


We did the same transition for CARM to 1/sqrt(trans) signals, as last night (elog 9793).  The only difference is that for CARM actuation, we were using a -1*MC in the output matrix, rather than +1's for both ETMs. 

We then had a look at the relative sign and gain between the ALS DARM signals, and AS55Q, using a calibration line in DARM.  Before doing so, we used the DARM line (521.3 Hz, 50 counts) to rotate the AS 55 phase from -60.7 degrees to -97.7 degrees, which gave us about 20dB separation between the I and Q signals.  This informed us that we needed a factor of about 400 less gain for AS55Q than for the ALS darm signal, as well as a minus sign, so I put -400 in the DC normalization place in the LSC for AS55, so that my input matrix would go from ALSY-ALSX (1's) to +1 in AS55Q. 

This transition to AS55 was very easy, and once we did it, we held lock for 5 or 10 minutes, until a large earthquake from Papa New Guinea hit us.  Note however, that we still had a large CARM offset, and our TRX and TRY signals were about 0.1 counts, when we expect several hundred at perfect resonance. 

After that, we relocked, made both CARM and DARM transitions again, and tried to look at a CARM calibration line to see if we see CARM information in any of the REFL RF signals.  We lost lock after a few minutes (so, not related to our calibration line), so we didn't finish, but it looks like REFL55I, normalized by TRX+TRY is the most promising.  Also, REFL55's phase was already very good, while REFL11's phase was not. 


There were some moderate changes to the LSC model that happened, and matching screen changes.  I put in a switch just before the input triggering place of the CARM servo.  This allows us to switch from the "regular" input matrix, and a CESAR signal.  The inputs to the CESAR block are sqrtinv(TRX), sqrtinv(TRY), ALSX, ALSY and the output of the CARM row of the input matrix (so that we can have dynamic normalization of the RF signals).  I have exposed all of these changes in the input matrix screens.

I also modified slightly the ALS watch scripts, to include CARM and DARM servo filter watching, so now we can use the actual CARM and DARM servos.  We should make restore configure scripts for these!


The 2 gps times for when we made the transition from ALS DARM to AS55 DARM were 1081238160 and 1081240217.  We want to go back tomorrow, and extract some nice time series.

Here's a spectrum though, of the difference in noise between DARM on ALS, and DARM on AS55.  The CARM was always on 1/sqrt(Trans) signals during these spectra.  We have an enormous gain in high frequency noise performance once we switch to the RF signal, which is great.

IRlocks.pdf

  9796   Fri Apr 11 01:02:07 2014 JenneUpdateLSCCARM and DARM both on IR signals!!!!!!!!!

[EricQ, Jenne]

We're still working, but I'm really excited, so here's our news:  We are currently holding the IFO on all IR signalsNo green, no ALS is being used at all!!!!  

PRCL and MICH in REFL33, CARM on 1/sqrt(trans), DARM on AS55 Q.

CARM actuating on MC2, DARM actuating +ETMY, -ETMX. 

CARM offset is 1.9 counts, TRX averages about .1 counts. At this offset, we are able to transition CARM from ALS to DC Transmission signals and DARM from ALS to AS55Q. 

onlyIRspectra.pdf

  9795   Thu Apr 10 16:09:29 2014 SteveUpdateVACRGA scan at 75% pumping speed

Quote:

 

 The loaner controller is swapped in. It has  520 Hz rotation speed.  This speed use to be 680 Hz with our old one.

 

Attachment 1: 520HzpumpingSpeed.png
520HzpumpingSpeed.png
  9794   Thu Apr 10 10:52:15 2014 manasaSummaryIOOMC2_TRANS path in WFS servo

Quote:

Quote:

I've also turned on the MC2 TRANS path to gather some data over the weekend on how well or bad it works. Please turn it off on Monday.

 MC2_TRANS path in WFS servo turned OFF.

[From yesterday]

The MC had not been stable lately with WFS drifting constantly. I checked the servo and found that the MC_TRANS path was still running. It turned out that the autolocker script enables the TRANS path in the locking process. I have turned the MC_TRANS path servo inputs OFF and now it is no more a part of the WFS servo.

P.S. Jenne fixed the PMC alignment mostly in pitch to bring it up to 0.81 from 0.77. But the temperature fluctuations have still not got us to the sweet spot for optimum PMC trans.

  9793   Thu Apr 10 01:56:05 2014 JenneUpdateLSCCARM transitioned to IR error signals!

[Jenne, EricQ]

This evening we took things a little bit farther than last night (elog 9791) and transitioned CARM to fully IR signals, no ALS at all for CARM error signals!  We were unsuccessful at doing the same for DARM. 

As we discussed at 40m Meeting this afternoon, the big key was to remove the PRCL ASC from the situation.  I don't know specifically yet if it's QPD saturation, or what, that was causing PRM to be pushed in pitch last night, but removing the ASC loops and reengaging the PRM optical lever worked like a dream. 

Since we can now, using ALS-only, get arbitrarily close to the PRMI+2 arm full resonance point, we decided to transition CARM over to the 1/sqrt(transmission) signals.  We have now done this transition 5 or 10 times.  It feels very procedural and robust now, which is awesome!

To make this transition easier, we made a proto-CESAR for the CARM signals in the LSC.  There's nothing automatic about it, it's just (for now) a different matrix. 


ALS lock conventions:

We have (finally listening to the suggestion that Koji has been making for years now....) set a convention for which side of the PSL the X and Y beatnotes should be, so that we don't have to guess-and-check the gain signs anymore.

For the X beatnote, when you increase the value on the slow slider, the beatnote should increase in frequency.  For the Y beatnote, when you increase the value on the slow slider, the beatnote should decrease in frequency. 

The input matrix (the aux input part) should then have +1 from ALSX->carm, and +1 from ALSY->carm.  It should also have -1 from ALSX->darm and +1 from ALSY->darm. 

The output matrix should be carm -> +1's for both ETMs.  darm should be -1 to ETMX and +1 to ETMY.

With these conventions, both carm and darm should have negative signs for their gains. 

Since we don't have (although should whip up) Watch scripts for the CARM and DARM servo filters, we were using the Xarm filterbank for carm, and the Yarm filterbank for darm again.


Transitioning CARM to 1/sqrt(trans) signals:

As with last night, we were able to easily acquire PRMI lock with a CARM offset of 3 counts.  We then moved down to 2 counts, and saw transmission values of 0.1-0.2.  We set the offsets in the TR_SQRTINV filter banks so that the difference between the outputs was zero, and the mean of the outputs was 2 (the same as the CARM offset we had). 

We looked at the relative gain and sign between the ALS and 1/sqrt() signals, and found that we needed a minus sign, and half the gain.  So, we stepped the 1/sqrt() matrix elements from 0 to -0.5 in steps of 0.1, and at the same time were stepping the ALS matrix elements to CARM from +1 to 0, in steps of 0.2.  This was, excitingly, very easy!

The first time we did this successfully, was a few seconds before 1081143556 gps.

Here is a set of spectra from the first time we locked on the 1/sqrt(trans) signals. 

 

sqrtInvLock.pdf

 


Failure to transition CARM to RF signals, or reduce CARM offset to zero:

While locked on the 1/sqrt(trans) signals, we looked at several RF signals as options for CARM.  The most promising seems to be REFL55, normalized by (TRX+TRY).  The next most promising looks like REFL11 normalized by POPDC.  Note that these are entirely empirical, and we aren't yet at the resonant point, so these may not be truly the best.  Anyhow, we need to reconfigure the LSC input of the normalized error signals, so that they can go into the CESAR matrices.  This was more than we were prepared to do during the nighttime.  However, it seems like we should be about ready to do the transition, once we have the software in place.  Right now, we either normalize both ALS and the RF signal, or we normalize neither.  We want to be able to apply normalization to only the RF signal. 

Just sitting on the tail of the CARM resonance, there were some random times when we seem to have swung through total resonance, and spoiled our 1/sqrt(trans) signals, which aren't valid at resonance, and so we lost lock.  This implies that auto-transitioning, as CESAR should do, will be helpful. 


Attempt at transitioning DARM to AS55:

Next up, we tried to transition DARM to AS55, after we had CARM on the 1/sqrt signals.  This was unsuccessful.  Part of the reason is that it's unclear what the relative gain should be between the ALS darm signals and AS55, since the transfer function is not flat.  Also, we didn't have much coherence between the ALS signals and AS55Q at low frequencies, below about 100 Hz, which is concerning.  Anyhow, more to investigate and think on here. 


Transitioning CARM to 1/sqrt signals, with a DARM offset:

As a last test, Q put in a DARM offset in the ALS control, rather than a CARM offset, and then was still able to transition CARM control to the 1/sqrt signals.  As we expect, when we're sitting on opposite sides of the arm resonances, the 1/sqrt signals have opposite signs, to make a CARM signal. 


Conclusions / path(s) forward:

We need to redo the LSC RF signal normalization, so that the normalized signals can be inputs to CESAR. 

We need to make sure we set the AS55 phase in a sane way.

We need to think about the non-flat transfer function (the shape was 1/f^n, where n was some number other than 0) between the ALS darm signal and AS55.  The shape was the same for AS55 I&Q, and didn't change when we changed the AS55 phase, so it's not just a phasing problem. 

What DC signals can we use for auto-transitioning between error signals for the big CARM CESAR?

  9792   Wed Apr 9 16:08:33 2014 JenneUpdateLSCCARM loop gains vs. CARM offset

I have taken EricQ's simulation results for the CARM plant change vs. CARM offset, and put that together with the CM and CARM digital control loops, to see what we have. 

The overall gains here aren't meaningful yet (I haven't set a UGF), but we can certainly look at the phases, and how the magnitude of the signals change with CARM offset.

First, the analog CM servo.  I use the servo shape from Den's elog from December, but only what he calls "BOOST", the regular servo shape, not any of the super boosts, "BOOST 1-3".   No normalization.

REFL11_analog.pngREFL55_analog.png

Next, the digital LSC CARM servo (same filters as XARM and YARM).  I have used FM4 and FM5, which are the 2 filters that we use to acquire regular LSC arm lock.  For the actuator, I just use a 1Hz pendulum as if I'm pushing only on the ETMs.

REFL11_digital.pngREFL55_digital.png

I also used the exact same setups as above, but normalized the transfer functions by a DC photodiode output.  The analog CM loops change the least (around a few kHz) if I use POPDC.  The digital CARM loops change the least (around 100Hz) if I use TRX (or, equivalently, TRX + TRY).

Here are the normalized plots:

REFL11_analog_normalizedPOPDC.pngREFL55_analog_normalizedPOPDC.png

REFL11_digital_normalizedTRX.pngREFL55_digital_normalizedTRX.png

Either way, with or without normalization, the digital CARM loop will go unstable between 0-10pm, for both the REFL RF photodiodes.  We need to figure out how to get a realistic transfer function out for the 1/sqrt(TRANS) signals, and see what happens with those.  If those also look unstable, then maybe we should consider a DC signal for the analog CM servo to start, since that could have a wider linear range.

  9791   Wed Apr 9 02:34:20 2014 JenneUpdateLSCJumping over the CARM resonance point

Koji was right, and I was using much too large of a CARM offset.  Tonight, I set either my CARM or DARM offset to 3 counts, and was able to easily acquire PRMI lock using REFL33. 

For either CARM or DARM offset reduction (the other one was kept at zero offset), I was able to get to about 0.5 counts, but I lose lock when I try to go to 0.4 or 0.3 counts.  One time, I tried "jumping over" the resonance, by going from minus 1 to plus 1 in CARM offset.  Plots of this below.


Locking notes

ALS locked with "Xarm" servo as proxy for DARM, and "Yarm" servo as proxy for DARM.  Pushing only on ETMs today, not the MC. 

MICH / PRCL:

Input matrix:  1's in REFL 33 I&Q, if not using power normalization.  200's in REFL 33 I&Q if power normalization used (either POPDC or POP22).  200 used because that's about the average value of POPDC or POP22 when PRMI sideband-only resonant.

Trigger:  POP22, up 100, down 10.

Power normalization:  1's for both MICH and PRCL in POP22I for one trial.  1's for both MICH and PRCL in POPDC for another trial.  Both seemed to work equally well, although that may change when I'm actually getting IR resonance in the cavity.

FM triggers:  MICH = FM2.  PRCL = FMs 2, 3, 6, 9.  Trigger up = 35, down 10.  PRCL delayed by 0.5 sec, MICH delayed by 5 seconds.

Servo gains:  MICH = 0.4, PRCL = -0.01


Observations:

When I approach the situation of both arms resonating, it pretty consistently looks like the PRM is getting pushed in pitch (and not in yaw).  I don't know why this could be, but it seems like this is the big symptom before lockloss - if the POP spot starts moving (and the PRM suspit signal starts moving), PRMI lock is going to be lost.

I don't know if it's imperfect alignment, imperfect mode matching, or something else, but I see lots of high-order higher order modes on both the POP and AS cameras when the CARM or DARM offset is less than 1 count.  I tried to take a video, but the brightness and contrast aren't set as high as on monitors 3 and 5, so it's hard to see the dim stuff.  Youtube.  At the midpoint of the video, you see a lockloss.

Even though I have overridden the transmission triggers so that I only use the QPDs for the transmission signals, I'm only seeing arm transmission values up to about 50 from each arm.  If we had ideal PRC gain, we expect something like 650 or 700. 


A few plots

All of the raw data for these plots, and several other channels, is in /users/jenne/PRFPMI/PRMI_2arms_8Apr2014/m1_to_p1_carmOffset_1081065069.  As mentioned above, "XARM" is CARM, and "YARM" is DARM.  So, the XARM_IN1 tells us about the CARM offset that I was applying.  The start time is 1081065069, and the plots are all 8 seconds long.

First, the transmitted power and the CARM offset.

TRX_TRY_QPDonly.png

The REFL_I error signals and the CARM offset.

LSC_error_signals.png

The RF signals that we will eventually chose from for CARM and DARM control. Note that I'm not sure about the AS55 phase, so I plot both I and Q.

REFL1f_AS55.png

The PRM suspit and sus yaw angular signals and the CARM offset.  I don't see a huge change in the suspit signal, but it does seem to change character once we approach arm resonances.

PRM_SUSangles.png

  9790   Tue Apr 8 23:04:14 2014 manasaUpdateLSCarm length measurements

Quote:

 Since we don't have an arm length precision measurement (i.e. better than centimeters), why not just do as Koji suggests and use the ALS to get the frequency spacing between a few red FSR and then we have the measurement solid ?

Arm lengths measured using ALS. Both the arms were estimated to have the same length (to the order of a centimeter) 37.51 m

How
I used ALS error signal to lock the arms and scanned the arm to find 4 consecutive IR resonances. From the beat note frequencies measured using the spectrum analyser during IR resonance, the FSR, and hence the length of the arms were calculated.

The estimated lengths are not very precise down to a mm given the resolution of the spectrum analyser. We have brought out the rubidium clock to use as reference for the spectrum analyser to challenge the measurements.

  9789   Tue Apr 8 19:10:39 2014 ranaUpdateLSCarm length measurements

 Since we don't have an arm length precision measurement (i.e. better than centimeters), why not just do as Koji suggests and use the ALS to get the frequency spacing between a few red FSR and then we have the measurement solid ?

  9788   Tue Apr 8 10:44:53 2014 KojiUpdateLSCPlaying with PRMI + 2 arms

I vaguely remember that the ALS count (Phase Tracker output) is converted to Hz@532nm by a factor  ~20kHz/cnt.
This means the calibration for the IR frequency is 10kHz/cnt.

If this is true 100cnt is 2MHz. Isn't it too big?

Assuming 38.5m for the arm length, FSR is 3.89MHz. (~389cnt)

Our sideband is at integer multiples of 11.03MHz. So...

1xf1 is 0.62MHz (62cnt) away from the carrier
2xf1 is 1.24MHz (124cnt)
3xf1 is 1.86MHz (186cnt)
5xf1=1xf2 is 0.79MHz (79cnt)
10xf1 = 2xf2 is 1.58MHz (158cnt)
15xf1 = 3xf2 is 1.52MHz (152cnt)

So we have to be well with in 62cnt to avoid resonating modulation sidebands.

There maybe some mistake in the factors.
e.g. Phase Tracker calibration is not correct, or CARM ALS OFFSET has factor 2 different calibration from the arm ALS offset.

  9787   Tue Apr 8 01:58:53 2014 JenneUpdateLSCPlaying with PRMI + 2 arms

I played around with the PRMI + 2 arms situation again this evening.  (I'm not ready to call it "PRFPMI" until we're at least partially using IR signals for CARM and DARM control).

I'm still a little bothered by the fact that we lose almost all light in the PRC when we're reducing the CARM offset.  I'm not sure where the light is going, but it's not circulating in the PRC, since I see the POP camera get very dark.  I can bring back light by changing either the PRCL offset, the MICH offset, or to a lesser extent (or maybe I'm not going far enough in offset) the DARM offset.

Tonight, I was using ALS to push on the ETMs in DARM/CARM mode (I didn't push on MC2 today, since it was being finicky and I had a hard time locking CARM with the MC as the actuator today). 

For the PRMI, I was using REFL 33 I&Q.  PRCL gain was -0.04, MICH gain was +0.8.  REFL 33I varied between +2 and +1.2 (smaller gain necessary as CARM offset was reduced, but it's easier to acquire at large CARM offset with larger gain), and REFL 33Q was +1.0.   PRCL has the usual FM 2,3,6,9 triggered, but MICH only had FM 2 triggered.  The others (particularly FM6) make MICH lose lock.  PRCL ASC was engaged, with the PRM oplev off.

Most of what I was trying tonight was to reduce the CARM offset, and then adjust some other offset (MICH, PRCL or DARM) to maximize POP DC.  It's possible that the POP 110 and 22 diodes are changing their optimal demod phases as the optical plant changes, but POPDC is just DC. 

I wasn't able to get the CARM offset below about 40 counts.  At some point I had both CARM and DARM offsets at 50 counts, and had IR resonance in the Xarm, but no resonance in the Yarm.  I guess this is just part of having a simultaneous CARM and DARM offset.

EDIT:  If I leave the CARM offset at 0, and use a large DARM offset (100 counts), I can acquire PRMI lock, but I run into the same problems as I reduce the DARM offset - the POP power decreases, and I lose lock around 45 counts.


Side note: Earlier today I redid the POP 110 and 22 phases.  POP110 was -61 deg, and is not -101 deg.  POP22 was +164 deg, and is now -105 deg.  I'm not sure why they needed such radical changes.  When sideband locked on REFL33, POP110 had an average DC level of 580 cts, while POP22 had a value of 290 cts.

  9786   Mon Apr 7 15:26:32 2014 jamie, ericqUpdateCDSaborted attempt to update c1sus machine with second CPU

This morning we attempted to replace the c1sus front end machine with a spare that had been given a second CPU, and therefore 6 additional cores (for a total of 12).  The idea was to give c1sus more cores so that we could split up c1rfm into two separate models that would not be running on the hairy edge of their cycle time allocation.  Well, after struggling to get it working we eventually aborted and put the old machine back in.

The problem was that the c1sus model was running erratically, frequently jumping up to 100 usec of a 60 usec clock allocation.  We eventually tracked the problem down to the fact that the CPUs in the new machine are of an inferior and slower model, than what's in the old c1sus machine.  The CPU were running about 30% slower, which was enough to bump c1sus, which nominally runs at ~51 usec, over it's limit.

This is of course stupid, and I take the blame.  I skimped on the CPUs when I bought the spare machines in an attempt to keep the cost down, and didn't forgot that I had done that when we started discussing using one of the spares as a c1sus replacement.

I think we can salvage things, though, by just purchasing a better CPU, one that matches what's currently in c1sus.  I'll get Steve on it:

c1sus CPU: Intel(R) Xeon(R) CPU X5680 3.33GHz

In any event, the old c1sus is back in place, and everything is back as it was.

Attachment 1: Screenshot-Untitled_Window.png
Screenshot-Untitled_Window.png
  9785   Fri Apr 4 18:51:29 2014 ericqSummaryLSCMORE CARM related modeling

In today's ISC call, Kiwamu was comparing two ways to approach resonance: 

  • "C-Type": The scheme we currently think about; zero DARM offset and slowly reduce the CARM offset
  • "D-Type": Start with no CARM offset, but a DARM offset and reduce that. 

D-type might be interesting to check out, since things change a little less dramatically when you reduce the DARM offset. Maybe this makes signal hopping easier? Signal recycling may complicate things, though. 

So, I've simulated CARM and DARM offset effects on CARM and DARM signals. (As with the previous plots, this is for the PRFPMI configuration.) From moving both offsets around, it looks like the resonance peak is about 5x wider in DARM than in CARM, so I simulated a 50pm offset range for CARM and a 250pm offset range for DARM. 

Here are some CARM signal transfer functions subject to CARM offsets in the top plot, and DARM offsets in the bottom plot. 

 carm2REFL11.pdfcarm2REFL55.pdf

carm2REFLDC.pdfcarm2TRX.pdf

 

It's looks like the DARM offset changes cause much less dramatic changes in the CARM plant features. It's conceivable that this would make CARM locking easier. 

Here are some DARM plant transfer functions. 

 darm2AS11.pdfdarm2AS55.pdf

darm2ASDC.pdfdarm2TRX.pdf

In these plots, I did something kind of artificial: when we move the CARM offset, it changes the proper demodulation phase to get DARM in the Q of the AS 1F RFPDS. So, at each CARM offset, I re-phased the AS 1F demodulators, to show the total DARM information available at the AS RFPDs at each offset, rather than what one would actually see in them with a static demod phase. 

ASDemodAngles.pdf 

  9784   Thu Apr 3 18:55:10 2014 ericqSummaryLSCSome CARM related modeling

 The other day, Jenne and I were comparing my MIST simulation to her Optickle simulation for the CARM transfer functions I posted some days ago. She told me that the arms are not exactly where they should be for the whole "PRC length tuning to account for sideband reflection phase off resonant cavity" deal. 

Specifically, as in the wiki (but with newer modulation frequencies), I calculated the ideal arm length to be 37.795 m some time ago, when doing PRC length simulations, and Jenne has told me that the X arm is more like 37.6m, and Y is 37.9. So, I updated my simulations, and found the following:

This does weird things to the f2 sideband buildup on resonance in the PRFPMI configuration:

asIsPRCRes.pdf idealPRCRes.pdf

(POP is way huger than than the TR's, because the POP pd's are artificially "inside" the cavity, whereas TRX/Y is actually transmitted through an ETM)

This is not necessarily directly something to worry about, but I think the following may be. It looks like this arm length mismatch actually causes the PRCL demodulation phase in REFL 165 to change dramatically with the CARM offset. (REFL33 seems fine, though. 5 degrees causes less than a 1% effective gain change.) 

 3fPrclAnglesCarm.pdf

My simulations don't include any signal recycling yet, so I don't have anything to show if there is a similar effect for SRCL, but it wouldn't surprise me... 

 

  9783   Thu Apr 3 18:09:54 2014 ericqUpdateLSCAttempt at PRMI+2 arms

So, here's the basic: "We reduced the CARM offset and saw more TRY" plot. 

offsetTRY.pdf

 

As Jenne mentioned, we suspected that we were seeing real displacement information in the sqrtInv signals. (We had incidentally hard switched to the transmon QPDs for all of this)

Here's a 2d-histogram of the ALS CARM error signal vs. the sqrtInv CARM signal (i.e. 1/sqrt(TRX) + 1/sqrt(TRY))

 

2dhist.pdf

This is exactly the shape we expect, which is cool. You can see where we stepped the offsets, too. It looks like the signal gets into it's good linear range when ALS CARM was about -20, which is when TRY was a little under 0.1, which seems pretty early and potentially useful. 

Also, here are snapshots of what REFL11_I and sqrtInv CARM were doing in the last five seconds of time in the above plot, which was shortly before we made the offset push that broke the PRMI lock. If you look really closely, maybe you can convince yourself that there is some common information in them...? It's hard to say. In any case, there is definitely CARM pdh action happening.

pdh_trans.pdf

  9782   Thu Apr 3 17:05:52 2014 SteveUpdateVACMaglev controller swapped

Quote:

Quote:

Quote:

 The date is correct on this monitor.

Last stored RGA scan data Dec 20, 2013

IFO pressure at CC1 5.8e-6 Torr

Valve configuration: Vacuum Normal, confirmable only by manual checking of position indicators and pressure gauge controllers  readouts

 

 The Osaka TG390MCAB maglev turbo pump's controller TC010M has passed the 40,000 hrs of operation. This triggered the " alarm" LED  warning light to come on. 

It is normal maintenance.  Maglev TP-1 is running perfectly.  Osaka will send us a loaner-controller that we can use while they do the std maintenance.

I'm thinking of ~ February to do this.

 We just received the loaner controller that will be swapped in it tomorrow morning.

The vacuum pressure will rise somewhat during this action.

 The loaner controller is swapped in. It has  520 Hz rotation speed.  This speed use to be 680 Hz with our old one.

Attachment 1: loanerControllerIn.png
loanerControllerIn.png
ELOG V3.1.3-